# justification=centering

### v1

@@ 6.3 Reliability of the contrast curves
Evaluating the adopted contrast curve metric through fake companion injections helps verify the consistency of its core assumptions, including the applied noise threshold, the definition of the median contrast curve, and the expected 50% recovery rate of injected signals. We inject fake point sources at the predicted sensitivity limit using the VIP package on a cluster center and examine their recovery across different RSM maps. Figure 12 compares the recovery of injected signals using RSM maps against the signal-to-noise ratio (S/N) map produced by PCA with five principal components, for the center of cluster 1. The flux levels of the injected companions have been chosen to be just above the expected RSM sensitivity limits as illustrated in Fig. 13, which also compares the detection limits of RSM with the full-frame PCA $5\sigma$ contrast curve. The injections are performed at multiple separations, with one injection at each separation and varying position angles. As expected, a significant fraction of the injections are successfully detected above the threshold in RSM maps, achieving approximately 50% recovery consistent with the definition of the contrast curve. Conversely, the same injections all appear well below the $5\sigma$ level in the S/N PCA map. This highlights RSM’s ability to deliver deeper detection limits and reliable contrast curves, surpassing conventional post-processing methods in HCI.
@@ 6.4 Improvement to the SHINE survey
Figure 14: Comparison of RSM and full-frame PCA contrast curves for the F150 SHINE sample: A visualization of all RSM contrast curves (shaded blue) and PCA contrast curves (shaded pink), including their median values, highlighting RSM’s significant improvement in detection limits at both small and large separations.
To conclude this analysis of sensitivity limits, we present the detection limits for the F150 SHINE sample derived from 213 observations using the detection threshold and modifications outlined in this study. Figure 14 shows all RSM contrast curves (shaded blue) against full-frame PCA ADI curves (shaded pink). For consistency within the survey, we do not use a high-pass filter at large separations, although we acknowledge that it could deliver better results in these cases. The PCA-ADI curves may show slight differences from those published in 2021A&A...651A..71L, reflecting variations in the data reduction process. Nevertheless, our PCA-ADI results align well with those reported in 2023A&A...675A.205C, as both are based on the same reduction pipeline. The PCA contrast curves display outliers due to bright companions and background star signals. The median contrast curves indicate RSM’s substantial improvement, achieving a factor 2 enhancement at 800 mas and a factor 4 to 5 at 135 mas compared to the five-component PCA results.
@@ 7 Identification of point sources
The application of the RSM algorithm to the F150 SHINE sample resulted in the identification of 87 signals across 59 different observations, while 154 observations from the F150 sample showed no detectable signal within the $1\aas@@fstack{\prime\prime}1$ field of view. As previously described, a signal is considered significant when it lies above the detection threshold in all four RSM detection maps, considering the two selected algorithm combinations and the parallel observations performed by SPHERE-IRDIS in the H2 and H3 filters. A dedicated tool was developed to cross-match detections between maps within half a FWHM, ensuring consistent identification of point sources across filter combinations. The photometric and astrometric measurements were derived following the procedure detailed in 2022A&A... [^6] A..33D. The performance of these measurements was assessed by (2024SPIE13097E..13C) during the Exoplanet Imaging Data Challenge (phase II); this analysis is not discussed further here, as it lies beyond the scope of the present work.
Among the detected signals, 49 were already recovered by 2025A&A...697A..99C, while 38 were not identified with PACO and required a dedicated analysis to distinguish genuine astrophysical sources from false positives. This counting of false positives does not include some artifacts that tend to appear in RSM maps in the presence of bright companions, at the same angular distance as those companions. This section presents a detailed comparison between the signals detected with RSM and PACO in Sect. 7.1, followed by an analysis of the new point sources recovered exclusively with RSM in Sect. 7.2 using proper motion and color-magnitude diagram tests.
@@ 7.1 Comparison with PACO on the F150 sample
Given the comparable performance of RSM and PACO, a direct comparison of their detected signals provides valuable insight into the relative sensitivity of both methods, the reliability of faint detections, and potential missed detections on either side.
@@
Out of the 52 PACO detections, 49 were also recovered by RSM based on our four-map detection criterion. Figure 17 presents the contrast values of each PACO-detected signal, with colors indicating the recovery status in RSM; dark blue points represent signals fully recovered in four maps, while other colors correspond to non-recovered sources. As shown in the figure, most signals detected by PACO were also recovered by RSM, with only a few exceptions.
@@
The non-recovered signals correspond to the confirmed planet HD 95086 b (in two observations), and a background star around HD 151726, although these sources were recovered in at least one of the RSM maps. The primary reason for this discrepancy lies in the different detection criteria employed by the two algorithms. The RSM method requires recovery above the detection threshold in both H2 and H3 filters, whereas PACO derives a single signal-to-noise ratio (S/N) that combines information from both spectral channels. Consequently, sources with marginal S/N in one band but a stronger signal in the other may be recovered by PACO but not by RSM.
This behavior is clearly illustrated by the case of HD 95086 b, observed on 2015-05-05 and 2015-05-11, with PACO S/N values of 6.8 and 5.0, respectively. 2022A&A...664A.139D showed that the planet/star contrast for this target is twice as large in H3 compared to H2 (see their Fig. 5), consistent with the 2025A&A...697A..99C database, where the H2 contrast errors are nearly as large as the measured contrast values themselves.
@@ 7.2 Identification of new point sources
The application of the RSM algorithm to the F150 SHINE sample revealed 38 newly detected point sources across 29 different observations for 27 stars. Among these targets, 10 stars have additional epochs within the F150 sample, while the remaining 17 correspond to single-epoch observations.
Multi-epoch data enable the differentiation between false positives and genuine astrophysical sources through proper motion tests.
Out of the ten targets that have additional epochs, two exhibit detections in more than one epoch, while the remaining eight show detections in only a single epoch despite the availability of multiple observations.
@@
Regarding the multi-epoch detections, RSM successfully recovered PZ Tel in an observation from 2014, which was not detected by PACO, unlike other epochs where the signal was identified by both methods. RSM also recovered a background star around HD 164249 identified through the proper motion test between the 2015 and 2016 epochs. These signals were not reported by 2025A&A...697A..99C but were previously mentioned by 2021A&A...651A..71L.
Figure 18: Color–magnitude diagram showing the new RSM detections in the F150 SHINE sample, plotted alongside the signals reported in 2025A&A...697A..99C for confirmed exoplanets, background stars identified through proper motion, and background stars classified via CMD analysis. The grey shaded area indicates the exclusion region for background stars defined by 2025A&A...697A..99C.
Verifying the nature of a signal for single-epoch targets is more challenging. In these cases, following 2025A&A...697A..99C, a color-magnitude diagram (CMD) analysis is employed to assess the likelihood of each detection being consistent with a planetary companion or a background star. The 17 targets that had no additional epochs within the SHINE F150 survey generated a total of 22 significant signals in our RSM maps. To these 22 signals, we add the ambiguous candidate identified in the previous paragraph, and perform a CMD analysis to evaluate the astrophysical relevance of each signal.
The CMD is shown in Fig. 18 for all the detections reported in 2025A&A...697A..99C, supplemented by the new RSM candidates. As discussed by 2025A&A...697A..99C, background stars occupy a well-defined region in the diagram, and we adopt their background-star region, referred to as the exclusion area in Fig. 18.
Most of the RSM candidates fall within this exclusion region, indicating that they are likely background stars. Two signals, falling outside the plot range of Fig. 18, display photometric properties inconsistent with an astrophysical source and are therefore classified as false positives. One last source lies just outside the exclusion region, exhibiting photometric characteristics consistent with an exoplanet-like nature, and is thus considered a promising planet candidate. Appendix I presents several examples of newly identified background-star detections and compares their corresponding PCA signal-to-noise ratios with the RSM scores.
@@
Figure 19: Summary of the detection statistics for the F150 sample, comparing the outcomes of RSM and PACO across four cases: no signal in either method, signal in PACO only, signal recovered by both methods, and signal recovered by RSM only. The dashed grey bars indicate the number of observations in each category, with the corresponding counts labeled above them. The colored stacked bars represent the number of detected signals, with colors denoting their classification: confirmed planets, background stars, ambiguous sources, false positives, and the newly identified promising candidate.
@@ 8 Conclusions
This study introduces new methods for determining detection thresholds in high-contrast imaging, utilizing the Regime Switching Model (RSM) applied to the SHINE F150 survey. Clustering based on simple environmental parameters using k-means effectively classified observations into meaningful groups, such as those with large number of frames, poor conditions, or specific effects like wind driven halo and low-wind effect. Spearman correlation tests confirmed the distinct noise characteristics in each cluster and validated the cluster representatives in terms of environmental and noise parameters. We have additionally highlighted how observational conditions influence RSM noise distribution in final RSM maps. By fitting a lognormal distribution at each separation, detection thresholds with a $3\times 10^{-7}$ FAP revealed notable variations across different clusters, emphasizing the importance of adapting thresholds to observational conditions. We also compared lognormal thresholds with those derived from maximizing the F1 score. While both approaches are compatible, the F1 score method better addresses individual bright false positives, making it suitable for single observations. Conversely, the lognormal fit reflects the overall noise behavior and provides a more generalizable solution for large surveys.
Leveraging the diverse noise behavior across frames in the science cubes and applying different PSF subtraction techniques further enhanced RSM’s ability to distinguish planetary signals from speckles. This led to narrower RSM noise distributions, reduced thresholds, and improved detection limits, achieving a fivefold improvement at 135 mas and a twofold overall enhancement at larger separations compared to 2021A&A...651A..71L for the F150 SHINE sample. Compared to 2025A&A...697A..99C, who used the PACO algorithm, applying RSM to the F150 sample achieved comparable performance. Our study led to the identification of 87 signals, 49 of which were previously reported by 2025A&A...697A..99C. Among the 38 new detections, 22 were classified as background stars, 14 as false positives, one turned out to be a known low-mass companion, and one was identified as a new exoplanet candidate for which follow-up observations are needed. This result highlights RSM’s robustness and potential for further applications.
Finally, the methodology developed in this work can be directly extended to the full SHINE survey and adapted to other ground-based high-contrast imaging programs, such as the Gemini-GPIES (2020AJ....159...71N), SPHERE-BEAST(2019A&A...626A..99J; 2021A&A...646A.164J), or NACO-ISPY (2020A&A...635A.162L) surveys. This framework provides a robust and versatile foundation for statistically consistent planet detection and performance characterization across a wide range of observing conditions.
@@
The RSM algorithm is publicly available on GitHub as a Python package [^1] https://github.com/chdahlqvist/RSMmap. The raw and reduced data used in this study can be requested through the High-Contrast Data Centre (HC-DC) [^2] https://hc-dc.cnrs.fr. The code developed for the analysis presented in this work is openly accessible on GitHub [^3] https://github.com/marisabalbal/RSM_f150SHINE. The full point-source candidate list will be presented in a forthcoming publication, together with an NA-SODINN analysis of the SHINE F150 sample (Cantero et al., in prep).

### v2

@@ 6.3 Reliability of the contrast curves
Evaluating the adopted contrast curve metric through fake companion injections helps verify the consistency of its core assumptions, including the applied noise threshold, the definition of the median contrast curve, and the expected 50% recovery rate of injected signals. We inject fake point sources at the predicted sensitivity limit using the VIP package on a cluster center and examine their recovery across different RSM maps. Figure 12 compares the recovery of injected signals using RSM maps against the signal-to-noise ratio (S/N) map produced by PCA with five principal components, for the center of cluster 1. The flux levels of the injected companions have been chosen to be just above the expected RSM sensitivity limits as illustrated in Fig. 13, which also compares the detection limits of RSM with the full-frame PCA $5\sigma$ contrast curve. The injections are performed at multiple separations, with one injection at each separation and varying position angles.
As expected, a significant fraction of the injections are successfully detected above the threshold in RSM maps, achieving approximately 50% recovery consistent with the definition of the contrast curve. Conversely, the same injections all appear well below the $5\sigma$ level in the S/N PCA map. This highlights RSM’s ability to deliver deeper detection limits and reliable contrast curves, surpassing conventional post-processing methods in HCI.
@@ 6.4 Improvement to the SHINE survey
Figure 14: Comparison of RSM and full-frame PCA contrast curves for the F150 SHINE sample. The illustration provides a visualization of all RSM contrast curves (shaded blue) and PCA contrast curves (shaded pink), including their median values, highlighting RSM’s significant improvement in detection limits at both small and large separations.
To conclude this analysis of sensitivity limits, we present the detection limits for the F150 SHINE sample derived from 213 observations using the detection threshold and modifications outlined in this study. Figure 14 shows all RSM contrast curves (shaded blue) against full-frame PCA ADI curves (shaded pink). For consistency within the survey, we do not use a high-pass filter at large separations, although we acknowledge that it could deliver better results in these cases. The PCA-ADI curves may show slight differences from those published in Langlois et al. (2021), reflecting variations in the data reduction process. Nevertheless, our PCA-ADI results align well with those reported in Chomez et al. (2023), as both are based on the same reduction pipeline. The PCA contrast curves display outliers due to bright companions and background star signals. The median contrast curves indicate RSM’s substantial improvement, achieving a factor of two enhancement at 800 mas and a factor of four to five at 135 mas compared to the five-component PCA results.
@@ 7 Identification of point sources
The application of the RSM algorithm to the F150 SHINE sample resulted in the identification of 87 signals across 59 different observations, while 154 observations from the F150 sample showed no detectable signal within the $1\aas@@fstack{\prime\prime}1$ field of view. As previously described, a signal is considered significant when it lies above the detection threshold in all four RSM detection maps, considering the two selected algorithm combinations and the parallel observations performed by SPHERE-IRDIS in the H2 and H3 filters. A dedicated tool was developed to cross-match detections between maps within half a FWHM, ensuring consistent identification of point sources across filter combinations. The photometric and astrometric measurements were derived following the procedure detailed in Dahlqvist et al. (2022). The performance of these measurements was assessed by (Cantalloube et al., 2024) during the Exoplanet Imaging Data Challenge (phase II); this analysis is not discussed further here, as it lies beyond the scope of the present work.
Among the detected signals, 49 were already recovered by Chomez et al. (2025), while 38 were not identified with PACO and required a dedicated analysis to distinguish genuine astrophysical sources from false positives. This counting of false positives does not include some artifacts that tend to appear in RSM maps in the presence of bright companions, at the same angular distance as those companions. This section presents a detailed comparison between the signals detected with RSM and PACO in Sect. 7.1, followed by an analysis of the new point sources recovered exclusively with RSM in Sect. 7.2 using proper motion and color-magnitude diagram tests.
@@ 7.1 Comparison with PACO on the F150 sample
Out of the 52 PACO detections, 49 were also recovered by RSM based on our four-map detection criterion. Figure 17 presents the contrast values of each PACO-detected signal, with colors indicating the recovery status in RSM; dark blue points represent signals fully recovered in four maps, while other colors correspond to unrecovered sources. As shown in the figure, most signals detected by PACO were also recovered by RSM, with only a few exceptions.
@@
The unrecovered signals correspond to the confirmed planet HD 95086 b (in two observations), and a background star around HD 151726, although these sources were recovered in at least one of the RSM maps. The primary reason for this discrepancy lies in the different detection criteria employed by the two algorithms. The RSM method requires recovery above the detection threshold in both H2 and H3 filters, whereas PACO derives a single signal-to-noise ratio (S/N) that combines information from both spectral channels. Consequently, sources with marginal S/N in one band but a stronger signal in the other may be recovered by PACO but not by RSM.
This behavior is clearly illustrated by the case of HD 95086 b, observed on 2015-05-05 and 2015-05-11, with PACO S/N values of 6.8 and 5.0, respectively. Desgrange et al. (2022) showed that the planet-star contrast for this target is twice as large in H3 compared to H2 (see their Fig. 5), consistent with the Chomez et al. (2025) database, where the H2 contrast errors are nearly as large as the measured contrast values themselves.
@@ 7.2 Identification of new point sources
The application of the RSM algorithm to the F150 SHINE sample revealed 38 newly detected point sources across 29 different observations for 27 stars. Among these targets, ten stars have additional epochs within the F150 sample, while the remaining 17 correspond to single-epoch observations.
Multi-epoch data enable the differentiation between false positives and genuine astrophysical sources through proper motion tests. Out of the ten targets that have additional epochs, two exhibit detections in more than one epoch, while the remaining eight show detections in only a single epoch despite the availability of multiple observations.
@@
Regarding the multi-epoch detections, RSM successfully recovered PZ Tel in an observation from 2014, which was not detected by PACO, unlike other epochs where the signal was identified by both methods. RSM also recovered a background star around HD 164249 identified through the proper motion test between the 2015 and 2016 epochs. These signals were not reported by Chomez et al. (2025) but were previously mentioned by Langlois et al. (2021).
Figure 18: Color–magnitude diagram showing the new RSM detections in the F150 SHINE sample, plotted alongside the signals reported in Chomez et al. (2025) for confirmed exoplanets, background stars identified through proper motion, and background stars classified via CMD analysis. The gray shaded area indicates the exclusion region for background stars defined by Chomez et al. (2025).
Verifying the nature of a signal for single-epoch targets is more challenging. In these cases, following Chomez et al. (2025), a color-magnitude diagram (CMD) analysis is employed to assess the likelihood of each detection being consistent with a planetary companion or a background star. The 17 targets that had no additional epochs within the SHINE F150 survey generated a total of 22 significant signals in our RSM maps. To these 22 signals, we add the ambiguous candidate identified in the previous paragraph, and perform a CMD analysis to evaluate the astrophysical relevance of each signal.
The CMD is shown in Fig. 18 for all the detections reported in Chomez et al. (2025), supplemented by the new RSM candidates. As discussed by Chomez et al. (2025), background stars occupy a well-defined region in the diagram, and we adopt their background-star region, referred to as the exclusion area in Fig. 18.
Most of the RSM candidates fall within this exclusion region, indicating that they are likely background stars. Two signals, falling outside the plot range of Fig. 18, display photometric properties inconsistent with an astrophysical source and are therefore classified as false positives. One last source lies just outside the exclusion region. It exhibits photometric characteristics consistent with an exoplanet-like nature and is thus considered a promising planet candidate. Appendix H presents several examples of newly identified background-star detections and compares their corresponding PCA signal-to-noise ratios with the RSM scores.
@@
Figure 19: Summary of the detection statistics for the F150 sample, comparing the outcomes of RSM and PACO across four cases: no signal in either method, signal in PACO only, signal recovered by both methods, and signal recovered by RSM only. The dashed gray bars indicate the number of observations in each category, with the corresponding counts labeled above them. The colored stacked bars represent the number of detected signals, with colors denoting their classification: confirmed planets, background stars, ambiguous sources, false positives, and the newly identified promising candidate.
@@ 8 Conclusions
This study introduces new methods for determining detection thresholds in high-contrast imaging, utilizing the RSM applied to the SHINE F150 survey. Clustering based on simple environmental parameters using k-means effectively classified observations into meaningful groups, such as those with large number of frames, poor conditions, or specific effects such as wind driven halo and low-wind effect. Spearman correlation tests confirmed the distinct noise characteristics in each cluster and validated the cluster representatives in terms of environmental and noise parameters. We have additionally highlighted how observational conditions influence RSM noise distribution in final RSM maps. By fitting a lognormal distribution at each separation, detection thresholds with a $3\times 10^{-7}$ FAP revealed notable variations across different clusters, emphasizing the importance of adapting thresholds to observational conditions. We also compared lognormal thresholds with those derived from maximizing the F1 score. While both approaches are compatible, the F1 score method better addresses individual bright false positives, making it suitable for single observations. Conversely, the lognormal fit reflects the overall noise behavior and provides a more generalizable solution for large surveys.
Leveraging the diverse noise behavior across frames in the science cubes and applying different PSF subtraction techniques further enhanced RSM’s ability to distinguish planetary signals from speckles. This led to narrower RSM noise distributions, reduced thresholds, and improved detection limits, achieving a fivefold improvement at 135 mas and a twofold overall enhancement at larger separations compared to Langlois et al. (2021) for the F150 SHINE sample. Compared to Chomez et al. (2025), who used the PACO algorithm, applying RSM to the F150 sample achieved comparable performance. Our study led to the identification of 87 signals, 49 of which were previously reported by Chomez et al. (2025). Among the 38 new detections, 22 were classified as background stars, 14 as false positives, one turned out to be a known low-mass companion, and one was identified as a new exoplanet candidate for which follow-up observations are needed. This result highlights RSM’s robustness and potential for further applications.
Finally, the methodology developed in this work can be directly extended to the full SHINE survey and adapted to other ground-based high-contrast imaging programs, such as the Gemini-GPIES (Nielsen et al., 2020), SPHERE-BEAST(Janson et al., 2019, 2021), or NACO-ISPY (Launhardt et al., 2020) surveys. This framework provides a robust and versatile foundation for statistically consistent planet detection and performance characterization across a wide range of observing conditions.
@@
The RSM algorithm is publicly available on GitHub as a Python package [^1] https://github.com/chdahlqvist/RSMmap. The raw and reduced data used in this study can be requested through the High-Contrast Data Centre (HC-DC) [^2] https://hc-dc.cnrs.fr. The code developed for the analysis presented in this work is openly accessible on GitHub [^3] https://github.com/marisabalbal/RSM_f150SHINE. The full point-source candidate list will be presented in a forthcoming publication, together with an NA-SODINN analysis of the SHINE F150 sample (Cantero et al., in prep). The table for the 213 F150 SHINE observations used in this study, along with their observing conditions, is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.